# L4P: Low-Level 4D Vision Perception Unified

### v2

@@ 5.2 Multi-Frame Optical Flow Estimation
We use the Spring dataset [37] for evaluation. We sample 289 16-frame clips from the train split. Spring is not used to train ours or other approaches we compare against, allowing us to evaluate generalization ability. The input frames are resized to $224\times 224$ for all evaluation. We use the Endpoint Error (EPE), as well as a more robust metric, ratio of EPE $<1$, for the evaluation.
We consider two baselines for comparison. RAFT [55], a competitive and widely used two-frame approach, creates dense pairwise pixel features and uses recurrent updates to estimate optical flow. MemFlow [10], a recently published work, ranks among the top methods on the Spring benchmark. It is a multi-frame approach that relies on a memory mechanism to leverage temporal context. Quantitatively, L4P compares favorably to both RAFT and MemFlow on Spring (Table 2). Our model captures both small and large motions and presents more precise motion boundaries (Figure 5, second row). In addition, multi-frame approaches like MemFlow and ours generally have an edge in temporal stability (see Supplementary). Unlike many specialized approaches, our model currently only operates on low-resolution videos and further work is needed to enable efficient high-res estimation.
@@
L4P $\mathbf{11.1}$ $\mathbf{17.7}$ $\mathbf{90.3}$ $6.0$ $10.5$ $\mathbf{87.8}$ $\mathbf{18.3}$ $\mathbf{28.0}$ $\mathbf{87.5}$ $\mathbf{11.8}$ $\mathbf{18.7}$ $\mathbf{88.5}$
@@
Figure 6: Sparse 2D/3D tracking on the TAPVid-3D benchmark. Comparison with SpaTracker, a SOTA 3D tracking approach. For joint visualization of depth and 3D tracks, we align them using median scaling. We use our depth maps for visualization of GT and for SpaTracker we use the ones used by their approach.
@@ 5.3 Sparse 2D/3D Track Estimation
We evaluate on TAPVid-3D [32], a benchmark containing around 2.1M long-range 3D point trajectories over three datasets: Aria [39], DriveTrack [53], and PStudio [25]. It introduced several baselines by combining SOTA 2D point tracking approaches, such as TAPIR [8], BootsTAPIR [9], and CoTracker [27], with depth solutions like ZoeDepth [2] and COLMAP [46, 48]. The top performing approach on the benchmark is SpaTracker [74].
@@
On 3D tracking, we outperform previous approaches on average across all the metrics (Table 4). Approaches that combine 2D track estimation with COLMAP perform better on the DriveTrack [53] dataset. A reason could be that many query points are on static vehicles, for which COLMAP gives accurate depth. Such COLMAP-based baselines, however, perform poorly on Aria [39] and PStudio [25], which are mostly dynamic. We show a comparison with SpaTracker (SOTA) in Figure 6.
@@ 5.5 Additional Tasks
Camera pose estimation. We follow [82] and represent a camera as a bundle of rays. More specifically, each camera is represented by the $16\times 16$ camera rays towards the image patch centers, and each ray by its 6-D Plücker coordinates. Our DPT dense head can be easily adapted to output a $16\times 16$ grid of rays. From the estimated ray bundles, we can then recover camera poses by solving least-square optimizations for camera centers and rotation matrices [82].
@@
L4P $\mathbf{0.115}$ $0.098$ $0.150$
For evaluation, we follow [81] and show results on both dynamic (Sintel and TUM-dynamics) and static scenes (ScanNet). We provide ATE (lower is better) and focus on feedforward approaches using general architectures. L4P performs favorably or on par with recent approaches like DUSt3R [65], Spann3R [60], CUT3R [64]. We show visual comparisons in Figure 5 (last row), and discuss additional metrics and methods in the Supplementary.
Discussion. Freezing the video encoder and fine-tuning a task-specific head is a simple way to add a new task. Better strategies may exist that allow for fine-tuning of the video encoder without affecting the performance of other tasks. Note that since we do not estimate in metric units, tasks like camera pose estimation and depth estimation can have different scales, albeit solved within a single unified framework. Adding constraints in the loss functions can help make the outputs of a new task consistent in scale with existing ones. We leave the investigation of these important problems to future work.
@@ 7 Acknowledgements
In this supplementary document, we provide comparisons for the inference time, and discuss additional implementation details regarding architecture, dataset and training. Please refer to supplementary video for a high-level overview and video results and comparisons.
@@ 1 Inference time
Our method solves multiple perception tasks at once and it is attractive for any application that requires a low-level, holistic understanding of the scene. We show that its performance is as good as, or better than methods designed for individual tasks. Its computational time also makes it appealing, as it solves these tasks jointly, with efficiency comparable with that of single-task methods. For a video clip of size $16\times 224\times 224$, our inference runs in around 300ms on an NVIDIA A6000, which corresponds to $\sim$19ms/frame, or $\sim$28ms/frame if we consider the overhead due to overlapping windows in long videos. This runtime includes all the tasks we discuss in the paper: depth, flow, tracking, motion segmentation, and pose estimation. Table 1 shows the runtime of the SOTA methods we compare against in the paper and confirms that our joint estimation is indeed efficient.
@@ 3 Architecture
For the sparse head, we adapt two-way attention and mask decoding from SAM [29]. We make appropriate modifications to make it work with video tokens and for our point tracking task. For an input point query, we use 3D positional encoding of dimension $C$ to generate a point query token. To identify all our input/output tokens we add learnable embeddings of dimension $C$. These input/output tokens interact with the input video tokens via a two-way attention mechanism. Following SAM, we only keep two layers of two-way attention mechanism. In the mask decoding stage, we add rearrangement operations to map video tokens of size $P\times C$ to video feature maps of size $T\times H\times W\times C/8$. We replace the 2D convolutions in the original mask-decoder of SAM with 3D convolutions. The video feature maps are finally mapped to 2D heatmap, depth-map and visibility map by doing a dot-product with the corresponding output tokens processed via the two-stage attention mechanism, as shown in Figure 3 in the main paper.
@@ 4 Training
For depth, we use SILog [11] loss, for flow, we use L1-loss on the estimated uv-offsets, and for tracking, we use L1-loss for the 2D track positions, scale-invariant loss for the track depth (similar to dense depth), and binary-cross entropy loss for the track visibility. We found the loss weights empirically by first weighting the losses to be in the same order of magnitude and then doing a small hyperparameter search around those weights. We use the loss weights of $20$ for the flow and depth losses. For tracking, we use loss weights of $1.0$, $20.0$, and $15.0$ for 2D track position loss, track depth loss and track visibility loss, respectively.
@@ 6 Camera pose estimation
DROID-SLAM∗† [56] $0.175$ $0.084$ $1.912$ $\mathbf{0.021}$ $\mathbf{0.013}$ $\mathbf{0.358}$ $\mathbf{0.046}$ $\mathbf{0.015}$ $\mathbf{0.449}$
@@
Spann3R [60] $0.329$ $0.110$ $4.471$ $0.056$ $0.021$ $0.591$ $\mathbf{0.096}$ $0.023$ $0.661$
CUT3R [64] $0.213$ $0.066$ $\mathbf{0.621}$ $\mathbf{0.046}$ $\mathbf{0.015}$ $\mathbf{0.473}$ $0.099$ $\mathbf{0.022}$ $\mathbf{0.600}$
L4P $\mathbf{0.115}$ $0.061$ $2.491$ $0.098$ $0.017$ $3.296$ $0.150$ $0.036$ $3.862$
L4P w/ intrinsics $\mathbf{0.115}$ $\mathbf{0.039}$ $0.844$ $0.099$ $\mathbf{0.015}$ $1.047$ $0.150$ $0.031$ $1.093$
Table 2: Additional camera pose estimation results. We consider two types of approaches: optimization-based and feedforward approaches. $\dagger$ indicates methods that require camera intrinsics at inference time, and * indicates the use of global optimization. “DUSt3R (w/o GA)” is an online variant of DUSt3R, which aligns all video frames with first frame, without using global alignment (GA).
@@
Training. We freeze the video encoder and train the camera pose estimation head with all datasets described in Section 2. GT rays are constructed for all samples using provided camera intrinsics and poses. We apply L2-loss on the Plücker coordinates, and use 16-frame clips for training. We use a batch size of 8 and train for 250k iterations.
Testing. At inference time, output Plücker coordinates can be converted to traditional camera pose representations for evaluation by solving least-square optimizations. Following [82], we consider and evaluate under two settings, one where camera intrinsics are provided at inference time, and one where they are not. The evaluation datasets, Sintel, TUM-dynamics, and ScanNet, all contain sequences longer than 16 frames, and therefore require inference beyond a single $T=16$ window. In order to do so, we perform inference on 16-frame windows with stride 8, and align the predictions using similarity transformation between consecutive windows. The similarity transformation is estimated from the overlapping frames to minimize alignment errors. We compare with both optimization-based and feedforward approaches in Table 2.

### v3

@@ 5.2 Multi-Frame Optical Flow Estimation
We use the Spring [42], Sintel [7], and Virtual KITTI [8] datasets for evaluation. These video datasets are not used to train ours or other approaches we compare against, allowing us to evaluate generalization ability. For each dataset, we use all the videos and all the non-overlapping 16-frame clips from each video. The input frames are resized to $224\times 224$ for all evaluation. We use the Endpoint Error (EPE), as well as a more robust metric, ratio of EPE <1<1, for the evaluation.
We consider two baselines for comparison. RAFT [60], a competitive and widely used two-frame approach, creates dense pairwise pixel features and uses recurrent updates to estimate optical flow. MemFlow [14], a recently published work, ranks among the top methods on the Spring benchmark. It is a multi-frame approach that relies on a memory mechanism to leverage temporal context. Quantitatively, L4P compares favorably to both RAFT and MemFlow (Table 2). Our model captures both small and large motions and presents more precise motion boundaries (Figure 4, second row). In addition, multi-frame approaches like MemFlow and ours generally have an edge in temporal stability (see Supplementary). Unlike many specialized approaches, our model currently only operates on low-resolution videos and further work is needed to enable efficient high-res estimation. We discuss this in the Limitations (Section 6).
@@
L4P $\mathbf{11.1}$ $\mathbf{17.7}$ $\mathbf{90.3}$ $6.5$ $11.3$ $\mathbf{87.8}$ $\mathbf{18.4}$ $\mathbf{28.1}$ $\mathbf{87.5}$ $\mathbf{12.0}$ $\mathbf{19.0}$ $\mathbf{88.5}$
@@
Figure 5: Comparisons on 2D/3D tracking (a). L4P can estimate camera poses, depth, and tracking in the same reference system (b).
@@ 5.3 Sparse 2D/3D Track Estimation
We evaluate on TAPVid-3D [35], a benchmark containing long-range 3D point trajectories over three datasets: Aria [44], DriveTrack [58], and PStudio [28]. It introduced several baselines by combining SOTA 2D point tracking approaches with depth solutions (eg. ZoeDepth [3], COLMAP [51, 53]). We mainly compare against CoTracker [30], TAPIR [12], and BootsTAPIR [13] for 2D tracking, and SpaTracker [80] for 3D tracking.
@@
On 3D tracking, we outperform SpaTracker (Table 4), demonstrating the effectiveness of our general architecture compared to a specialized one. Approaches that combine 2D track estimation with COLMAP perform better on the DriveTrack [58] dataset. A reason could be that many query points are on static vehicles, for which COLMAP gives accurate depth. Such COLMAP-based baselines, however, perform poorly on Aria [44] and PStudio [28], which are mostly dynamic. We show a comparison with SpaTracker in Figure 5.
@@ 5.5 Additional Tasks
Camera pose estimation. We follow [90] and represent a camera as a bundle of rays. More specifically, each camera is represented by the $16\times 16$ camera rays towards the image patch centers, and each ray by its 6-D Plücker coordinates. Our DPT dense head can be easily adapted to output a $16\times 16$ grid of rays. From the estimated ray bundles, we can then recover camera poses by solving least-square optimizations for camera centers, rotation matrices and even camera intrinsics [90].
@@
L4P (head FT) $0.162$ $0.060$ $0.100$
L4P (e2e FT) $\mathbf{0.132}$ $0.055$ $\mathbf{0.080}$
For evaluation, we follow [89] and show results on both dynamic (Sintel and TUM-dynamics) and static scenes (ScanNet). We provide ATE (lower is better) and focus on feedforward approaches using general architectures. We show two versions: one by only fine-tuning the camera head (head FT), and one by fine-tuning the entire model end-to-end (e2e FT) for all tasks, including camera pose estimation. In either settings, our approach performs favorably with recent approaches like DUSt3R [71], Spann3R [65], CUT3R [70]. Once we have the camera pose estimation, we can visualize the camera trajectory, depth and 3D tracks in the same reference system (Figure 5(b)). Additional metrics and methods are in the Supplementary.
6 Limitations
Our main limitation is the low input resolution ($224\times 224$), inherited from VideoMAE to leverage pretrained weights. This limits the use of high-resolution information useful for many applications. We emphasize this is not a fundamental limitation, and could be addressed through high-resolution fine-tuning of VideoMAE (following DINOv2 [43]) or by incorporating convex upsampling layers (common in opticalflow approaches). A second limitation is weaker performance on single-image depth estimation (Table 1, NYUv2), since L4P is trained only on videos. Incorporating image-only training into pre-training of VideoMAE and fine-tuning of L4P could help close this gap.
@@ 8 Acknowledgements
In this supplementary document, we provide comparisons for the inference time, and discuss additional implementation details regarding architecture, dataset and training. We also provide additional results for 4D reconstruction and 3D tracking. Please refer to supplementary video for a high-level overview and video results and comparisons.
@@ 1 Inference time
Our method solves multiple perception tasks at once and it is attractive for any application that requires a low-level, holistic understanding of the scene. We show that its performance is competitive with, and sometimes better than methods designed for individual tasks. Its computational time also makes it appealing, as it solves these tasks jointly, with efficiency comparable with that of single-task methods. For a video clip of size $16\times 224\times 224$, our inference runs in around 300ms on an NVIDIA A6000, which corresponds to $\sim$19ms/frame, or $\sim$28ms/frame if we consider the overhead due to overlapping windows in long videos. This runtime includes all the tasks we discuss in the paper: depth, flow, tracking, motion segmentation, and pose estimation. Table 1 shows the runtime of the SOTA methods we compare against in the paper and confirms that our joint estimation is indeed efficient.
@@ 3 Architecture
For the sparse head, we adapt two-way attention and mask decoding from SAM [32]. We make appropriate modifications to make it work with video tokens and for our tracking task. For an input point query, we use 3D positional encoding of dimension $C$ to generate a point query token. To identify all our input/output tokens we add learnable embeddings of dimension $C$. These input/output tokens interact with the input video tokens via a two-way attention mechanism. Following SAM, we only keep two layers of two-way attention mechanism. In the mask decoding stage, we add rearrangement operations to map video tokens of size $P\times C$ to video feature maps of size $T\times H\times W\times C/8$. We replace the 2D convolutions in the mask-decoder of SAM with 3D convolutions. The video feature maps are finally mapped to 2D heatmap, depth-map and visibility map by doing a dot-product with the corresponding output tokens processed via the two-stage attention mechanism, as shown in Figure 3 in the main paper.
@@
Figure 6: Scale consistency. To make the scales of different tasks consistent, we compute the scale factor from the depth loss and use it to constrain the scales for the 3D track depth and camera pose losses.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: 4D Reconstruction Results. We show example visualizations on diverse real-world sequences. For each scene, we jointly display the estimated depth maps, camera trajectories, and 3D point tracks, all aligned in a common reference frame. We also show the first and last input RGB frames in the top left corner for each example. Our approach enables efficient feedforward estimation of consistent and coherent 4D reconstructions across challenging dynamic environments.
@@ 4 Training
Losses. For depth, we use SILog [15] loss, for flow, we use L1-loss on the estimated uv-offsets, and for tracking, we use L1-loss for the 2D track positions, scale-invariant loss for the track depth (similar to dense depth), and binary-cross entropy loss for the track visibility. For the additional tasks of motion-based segmentation and camera pose estimation, we use binary cross-entropy loss and L1 loss, respectively. We found the loss weights empirically by first weighting the losses to be in the same order of magnitude and then doing a small hyperparameter search around those weights. We use the loss weights of $20$ for the flow and depth losses. For tracking, we use loss weights of $1.0$, $20.0$, and $15.0$ for 2D track position loss, track depth loss and track visibility loss, respectively.
The scale-invariant loss [15] we use for depth and 3D track depth is computed as shown below:
$L(y,y^{*})=\frac{1}{N}\sum_{i=1}^{N}\left(\log(y_{i})-\log(y_{i}^{*})+\alpha(y,y^{*})\right)^{2}$ (1)
$\alpha(y,y^{*})=\frac{1}{N}\sum_{i=1}^{N}\left(\log(y^{*})-\log(y)\right)$ (2)
where $y$ and $y^{*}$ are the predicted and ground-truth entities, respectively. For any prediction $y$, $e^{\alpha(y,y^{*})}$ is the scale that best aligns it to the ground truth.
Scale consistency. While we solve depth, 3D tracking and camera pose estimation within a single unified framework, they could have different scale factors if proper care is not taken. We need additional constraints in our loss functions to make them consistent with one another. To achieve this, we follow a simple approach and constrain the scale for the 3D track depth and camera pose losses to be the same as that of the depth loss. This is done by computing losses in a specific order as shown in Figure 6. First, we compute the scale-invariant loss and the scale factor $\alpha$ for the depth estimation following Equations 1 and 2. We then use the estimated scale factor from depth loss to constrain the scales for the 3D track depth loss and camera pose loss. This allows us to constrain the scales of different tasks and to visualize them together in a common reference frame, as shown in the 4D reconstruction visualizations in Figure 7.
@@ 6 Camera pose estimation
[…]
@@
Spann3R [65] $0.329$ $0.110$ $4.471$ $0.056$ $0.021$ $0.591$ $0.096$ $0.023$ $0.661$
CUT3R [70] $0.213$ $0.066$ $0.621$ $\mathbf{0.046}$ $\mathbf{0.015}$ $0.473$ $0.099$ $\mathbf{0.022}$ $0.600$
L4P (head FT) $0.162$ $0.072$ $0.999$ $0.060$ $0.019$ $0.935$ $0.100$ $0.031$ $0.940$
L4P (head FT) $0.149$ $0.072$ $0.417$ $0.062$ $0.017$ $0.552$ $0.095$ $0.029$ $0.584$
L4P (e2e FT) $0.132$ $0.056$ $0.577$ $0.055$ $0.019$ $0.675$ $0.080$ $0.028$ $0.705$
L4P (e2e FT) $\mathbf{0.115}$ $\mathbf{0.049}$ $\mathbf{0.379}$ $0.048$ $\mathbf{0.015}$ $\mathbf{0.462}$ $\mathbf{0.075}$ $0.025$ $\mathbf{0.477}$
Table 2: Additional camera pose estimation results. We consider two types of approaches: optimization-based and feedforward approaches. “DUSt3R (w/o GA)” is an online variant of DUSt3R, which aligns all video frames with first frame, without using global alignment (GA). $\dagger$ indicates methods that use camera intrinsics at inference time, and * indicates the use of global optimization. Our Plücker-based approach works with or without camera intrinsics. We show results for two versions. The “head FT” version only fine-tunes the camera head while keeping the L4P video encoder frozen. The “e2e FT” version fine-tunes the entire model end-to-end for all tasks (depth, flow, 2D/3D tracking and camera pose).
@@
Training. We freeze the video encoder and train the camera pose estimation head with all datasets described in Section 2. GT rays are constructed for all samples using provided camera intrinsics and poses. We scale the pose using the scale factor estimated from the depth loss to make the pose and depth scales consistent with each other. We apply L1-loss on the Plücker coordinates, and use 16-frame clips for training. We use a batch size of 8 and train for 200k iterations. We also train a version of the model that is trained end-to-end for depth, flow, 2D and 3D tracking, and camera pose estimation. We report the numbers for both head-only fine-tuning and end-to-end fine-tuning in Table 2.
Testing. At inference time, output Plücker coordinates can be converted to traditional camera pose representations for evaluation by solving least-square optimizations, as done in [90]. We consider and evaluate under two settings, one where camera intrinsics are provided at inference time, and one where they are not. The evaluation datasets, Sintel, TUM-dynamics, and ScanNet, all contain sequences longer than 16 frames, and therefore require inference beyond a single $T=16$ window. To handle sequences longer than 16 frames, we perform inference on overlapping 16-frame windows with a stride of 8. We then align the predictions between consecutive windows using a 3D similarity transformation. This transformation is computed using the estimated depths and camera poses from the overlapping frames of adjacent windows, leveraging the fact that our depth and camera pose predictions are trained to have a consistent scale. Specifically, we lift the depth maps to 3D point clouds using the estimated camera poses within each window, and then estimate the 3D similarity transform to align the camera poses across windows. We compare our results with both optimization-based and feedforward approaches, as shown in Table 2.
7 Additional results
4D reconstruction. We show additional visualization results for 4D reconstruction, where we jointly visualize the estimated depth maps and 3D point tracks aligned in a common reference frame using the estimated camera poses and camera intrinsics. These results are generated using our efficient feedforward approach. We show additional video results for 4D reconstruction in our supplementary video.
Additional 3D tracking results. Evaluating 3D tracking performance on the TapVid3D [35] benchmark is computationally intensive, given the large number of videos and annotations in the dataset. Therefore, in addition to the full evaluation results presented in the main paper, we also report our results on the minival subset, which contains 150 videos. The 2D AJ, 2D APD, 3D AJ, 3D APD and OA metrics for our approach are $49.0$, $65.0$, $10.9$, $17.5$ and $87.6$ respectively. We can compare these numbers with the minival results reported for Seurat [10], a recent 3D tracking method that utilizes 2D point tracks from the state-of-the-art CoTracker [30] and depth maps from DepthPro [6]. The 3D AJ and 3D APD numbers reported by authors are $11.4$ and $18.1$, which are slightly better than ours.